# ZTF-EP-25-07

### v1

@@ 6.1 First family
where $\alpha$ and $\beta$ are constants, and $\psi$ is an auxiliary function which will be defined below. As a “special ingredient”, Lagrangian (30) contains the boundary $\mathcal{G}$-term (in the action we have $\mathcal{G}\,\bm{\epsilon}=-2\,\mathrm{d}(\mathbf{A}\wedge\mathbf{F})$ with the volume 4-form $\bm{\epsilon}$ and the gauge 1-form $\mathbf{A}$), which does not affect the field equations, but it contributes to the charges, akin to the so-called Witten effect [Wit79, G+04]. Magnetic field is already fixed, $B_{r}=P/r^{2}$, while for the electric field we choose the ansatz [^4] This form of the electric field is reminiscent of the ModMax case [FAGMLM21, BH25a, BH25b], in which $E_{r}=Qe^{-\gamma}/r^{2}=(Q/P)e^{-\gamma}B_{r}$, although the ModMax Lagrangian doesn’t belong to the family (30). $E_{r}=\alpha B_{r}$. Consequently, we have $\mathcal{F}=2(1-\alpha^{2})P^{2}/r^{4}$ and $\mathcal{G}=4\alpha P^{2}/r^{4}$, so that $\mathcal{F}+\alpha\mathcal{G}=2(1+\alpha^{2})P^{2}/r^{4}$.
Now, given that the charges are related via $Q=4\beta P$, source-free NLE Maxwell’s equations are satisfied. Furthermore, first gravitational field equation (16) is reduced to
@@
where asterisk “*” denotes the substitution $r=(2P^{2}/x)^{1/4}$. Finally, using the first gravitational field equation and its derivative with respect to the radial coordinate $r$, it is straightforward to check that the second gravitational field equation is satisfied.
@@
Constant $\psi(0)$ is irrelevant, as it represents the constant term in the Lagrangian. Part of the weak field limit may be fixed if $\psi^{\prime}(0)=-(1+\alpha^{2})/4$, while $\mathcal{G}$-term in the Lagrangian may be effectively “removed” at least in the weak field limit with the choice $\beta=-\alpha\psi^{\prime}(0)/(1+\alpha^{2})=\alpha/4$. More concretely, if the metric function has an asymptotic form
@@ 6.2 Second family
Yet another family of regular black hole solutions is built from the Lagrangian ansatz with separated $\mathcal{F}$-part and $\mathcal{G}$-part [BFJS24],
@@
such that $\mathcal{K}(0)=0$ and $\kappa\mathrel{\mathop{:}}=\mathcal{K}_{\mathcal{G}}(0)\neq 0$. Here we choose a trivial electric field, $E_{r}=0$, and impose a constraint on charges $Q=4\kappa P$, which implies that NLE Maxwell’s equations are satisfied. Rest of the field equations
@@
is equivalent to the well-known magnetic case with the $\mathcal{F}$-dependent Lagrangian. Thus, using the usual Lagrangian reverse engineering (again, see Appendix A) we may choose a metric function $f(r)$ such that the total Lagrangian $\mathscr{L}(\mathcal{F},\mathcal{G})$ satisfies the qMWF limit (an obstacle for the full MWF limit is nonvanishing parameter $\kappa\neq 0$, which is essential for the presence of the electric charge in this construction). Note that, due to theorem 3 in [BSJ22a], $\mathcal{J}(\mathcal{F})$ cannot be simply Maxwell’s $\mathcal{F}/4$. In conclusion, we have a dyonic regular black hole, alas with the vanishing electric field. Overlap between two introduced families of Lagrangians and solutions occurs when $\mathcal{K}(\mathcal{G})=\kappa\mathcal{G}$, $\alpha=0$, and $\beta=\kappa$.
@@ 7 Discussion
Starting with the NLE Lagrangians of the form $\mathscr{L}(\mathcal{F},\mathcal{G})$, previous results [Bro01, BSJ22a], as well as novel theorem 5.1, have demonstrated that there are no static, spherically symmetric, electrically charged, regular black holes with $\mathscr{L}(\mathcal{F},\mathcal{G})$ Lagrangian respecting the qMWF limit. In other words, in order to obtain regular black hole solutions, one must admit at least one of the two features: irregular Lagrangian or magnetic charges. Former choice represents a theoretical challenge, as we need to make some sense of the weak field limit and connection with the canonical Maxwell’s electromagnetism. On the other hand, the introduction of magnetic charges, at least initially, doesn’t seem like a radical idea. It is true that at the moment of writing this paper, there is still no experimental evidence of magnetic charges in our universe. Nevertheless, there are some theoretical suggestions that they might exists and it is possible, in principle, that we need just a minuscule amount of magnetic charge in order to regularize the black holes.
Unfortunately, here we encounter a much serious obstacle. Analysis in section 4 reveals that a bounded Kretschmann scalar $K(r)$ necessitates that Komar mass $M$, electric charge $Q$ and magnetic charge $P$ cannot remain independent black hole parameters. Relation between mass and charges is not completely unheard of in theoretical physics, with examples such as dyonic BPS states [SW94]. However, given that we strive to give an argument for the regularity of generic static black holes, we have to reach for additional “theoretical gymnastics”. First of all, one needs to fine-tune the mass-charge relation in order to make it consistent with numerous astrophysical observations of macroscopic objects (we are not aware of any physical mechanism that would naturally lead to the desired black hole mass-charge constraint). Even worse, the implementation of the universal mass-charge relation at the level of known particles is bound to fail. Let us assume that the hypothetical magnetic charge is conserved and antiparticles carry opposite magnetic charge with respect to their particle counterparts. Now, both $\pi^{0}$ meson and $Z$ boson may decay into particle-antiparticle pairs, hence we may deduce that $P_{\pi^{0}}=P_{Z}=0$. Furthermore, $\pi^{0}$ and $Z$ are also electrically neutral, so that universal relation of the form $M=\mu(Q,P)$ would imply that $M_{\pi^{0}}=M_{Z}=\mu(0,0)$, which is certainly not true! Of course, none of these arguments do not completely eliminate the possibility of such a NLE theory, as one could extend the hypothetical mass-charge relation into $M=\mu(Q,P,\dots)$ with some additional variables (quantum numbers). Nonetheless, the obtained constraints inevitably push the construction of such a theory into the realm of, to put it mildly, unorthodox theoretical assumptions.
@@ Appendix A Regular black holes with rational metric function
with the abbreviation $\omega\mathrel{\mathop{:}}=(1+\alpha^{2})/(2^{1/2}\gamma)$. As can be easily verified, obtained function $\psi$ satisfies $\psi(x)=-(1+\alpha^{2})x/4+O(x^{5/4})$ as $x\to 0^{+}$. One of the basic assumptions $b_{3}\neq 0$ implies $\lambda_{3}\neq 0$, hence the denominator in (45) is nonzero at least on some interval $\left[0,x_{0}\right$>. As a consequence, function $\psi$, which is initially formally defined on the interval $\left[0,x_{0}\right$>, can be extended into a $C^{1}$ function across the origin $x=0$ via $\psi(-x)=-\psi(x)$ for $x\in\left<0,x_{0}\right$>.

### v2

@@ 6.1 First family
where $\alpha$ and $\beta$ are constants, and $\psi$ is an auxiliary function which will be defined below. As a “special ingredient”, Lagrangian (30) contains the boundary $\mathcal{G}$-term (in the action we have $\mathcal{G}\,\bm{\epsilon}=-2\,\mathrm{d}(\mathbf{A}\wedge\mathbf{F})$ with the volume 4-form $\bm{\epsilon}$ and the gauge 1-form $\mathbf{A}$), which does not affect the field equations, but it contributes to the charges, akin to the so-called Witten effect [Wit79, G+04]. Magnetic field is already fixed, $B_{r}=P/r^{2}$, while for the electric field we choose the ansatz [^5] This form of the electric field is reminiscent of the ModMax case [FAGMLM21, BH25a, BH25b], in which $E_{r}=Qe^{-\gamma}/r^{2}=(Q/P)e^{-\gamma}B_{r}$, although the ModMax Lagrangian does not belong to the family (30). $E_{r}=\alpha B_{r}$. Consequently, we have $\mathcal{F}=2(1-\alpha^{2})P^{2}/r^{4}$ and $\mathcal{G}=4\alpha P^{2}/r^{4}$, so that $\mathcal{F}+\alpha\mathcal{G}=2(1+\alpha^{2})P^{2}/r^{4}$.
Now, given that the charges are related via $Q=4\beta P$, source-free NLE Maxwell’s equations are satisfied. Furthermore, the first gravitational field equation (16) is reduced to
@@
where the asterisk “*” denotes the substitution $r=(2P^{2}/x)^{1/4}$. Finally, using the first gravitational field equation and its derivative with respect to the radial coordinate $r$, it is straightforward to check that the second gravitational field equation is satisfied.
@@
Constant $\psi(0)$ is irrelevant, as it represents the constant term in the Lagrangian. Part of the weak field limit may be fixed if $\psi^{\prime}(0)=-(1+\alpha^{2})/4$, while the $\mathcal{G}$-term in the Lagrangian may be effectively “removed” at least in the weak field limit with the choice $\beta=-\alpha\psi^{\prime}(0)/(1+\alpha^{2})=\alpha/4$. More concretely, if the metric function has an asymptotic form
@@ 6.2 Second family
Another family of regular black hole solutions is built from the Lagrangian ansatz with separated $\mathcal{F}$ part and $\mathcal{G}$ part [BFJS24],
@@
such that $\mathcal{K}(0)=0$ and $\kappa\mathrel{\mathop{:}}=\mathcal{K}_{\mathcal{G}}(0)\neq 0$. Here we choose a trivial electric field, $E_{r}=0$, and impose a constraint on charges $Q=4\kappa P$, which implies that NLE Maxwell’s equations are satisfied. The rest of the field equations
@@
are equivalent to the well-known magnetic case with the $\mathcal{F}$-dependent Lagrangian. Thus, using the usual Lagrangian reverse engineering (again, see Appendix A) we may choose a metric function $f(r)$ such that the total Lagrangian $\mathscr{L}(\mathcal{F},\mathcal{G})$ satisfies the qMWF limit (an obstacle for the full MWF limit is nonvanishing parameter $\kappa\neq 0$, which is essential for the presence of the electric charge in this construction). Note that, due to Theorem 3 in [BSJ22a], $\mathcal{J}(\mathcal{F})$ cannot be simply Maxwell’s $\mathcal{F}/4$. In conclusion, we have a dyonic regular black hole, alas with the vanishing electric field. Overlap between two introduced families of Lagrangians and solutions occurs when $\mathcal{K}(\mathcal{G})=\kappa\mathcal{G}$, $\alpha=0$, and $\beta=\kappa$.
@@ 7 Discussion
Starting with the NLE Lagrangians of the form $\mathscr{L}(\mathcal{F},\mathcal{G})$, previous results [Bro01, BSJ22a], as well as novel Theorem 5.1, have demonstrated that there are no static, spherically symmetric, electrically charged, regular black holes with $\mathscr{L}(\mathcal{F},\mathcal{G})$ Lagrangian respecting the qMWF limit. In other words, in order to obtain regular black hole solutions, one must admit at least one of the two features: irregular Lagrangian or magnetic charges. The former choice represents a theoretical challenge, as we need to make some sense of the weak field limit and connection with the canonical Maxwell’s electromagnetism. On the other hand, the introduction of magnetic charges, at least initially, does not seem like a radical idea. It is true that at the moment of writing this paper, there is still no experimental evidence of magnetic charges in our universe. Nevertheless, there are some theoretical suggestions that they might exist and it is possible, in principle, that we need just a minuscule amount of magnetic charge in order to regularize the black holes.
Unfortunately, here we encounter a much serious obstacle. Analysis in section 4 reveals that a bounded Kretschmann scalar $K(r)$ necessitates that Komar mass $M$, electric charge $Q$ and magnetic charge $P$ cannot remain independent black hole parameters. The relation between mass and charges is not completely unheard of in theoretical physics, with examples such as dyonic BPS states [SW94]. However, given that we strive to give an argument for the regularity of generic static black holes, we have to reach for additional “theoretical gymnastics”. First of all, one needs to fine-tune the mass-charge relation in order to make it consistent with numerous astrophysical observations of macroscopic objects (we are not aware of any physical mechanism that would naturally lead to the desired black hole mass-charge constraint). Even worse, the implementation of the universal mass-charge relation at the level of known particles is bound to fail. Let us assume that the hypothetical magnetic charge is conserved and antiparticles carry opposite magnetic charge with respect to their particle counterparts. Now, both the $\pi^{0}$ meson and $Z$ boson may decay into particle-antiparticle pairs, hence we may deduce that $P_{\pi^{0}}=P_{Z}=0$. Furthermore, $\pi^{0}$ and $Z$ are also electrically neutral, so that universal relation of the form $M=\mu(Q,P)$ would imply that $M_{\pi^{0}}=M_{Z}=\mu(0,0)$, which is certainly not true! Of course, none of these arguments do not completely eliminate the possibility of such a NLE theory, as one could extend the hypothetical mass-charge relation into $M=\mu(Q,P,\dots)$ with some additional variables (quantum numbers). Nonetheless, the obtained constraints inevitably push the construction of such a theory into the realm of, to put it mildly, unorthodox theoretical assumptions.
@@ Appendix A Regular black holes with rational metric function
with the abbreviation $\omega\mathrel{\mathop{:}}=(1+\alpha^{2})/(2^{1/2}\gamma)$. As can be easily verified, the obtained function $\psi$ satisfies $\psi(x)=-(1+\alpha^{2})x/4+O(x^{5/4})$ as $x\to 0^{+}$. One of the basic assumptions $b_{3}\neq 0$ implies $\lambda_{3}\neq 0$, hence the denominator in (45) is nonzero at least on some interval $\left[0,x_{0}\right$>. As a consequence, function $\psi$, which is initially formally defined on the interval $\left[0,x_{0}\right$>, can be extended into a $C^{1}$ function across the origin $x=0$ via $\psi(-x)=-\psi(x)$ for $x\in\left<0,x_{0}\right$>.